# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v1

@@ Introduction
   -w 5sec Audio Cond. 0.0111 0.0716 0.0021 0.0061 0.0058 0.0386 0.0605 0.0628 0.0124 0.0387
   -w 15sec Audio Cond. 0.0117 0.0750 0.0020 0.0064 0.0063 0.0320 0.0552 0.0525 0.0079 0.0398
@@
Table 1: Objective evaluation of EQ attribute improvement with average absolute error values – the lower, the better.
@@
   -w 5sec Audio Cond. 0.3717 0.3658 0.3919 0.3079 0.0164 0.0893 1.779 0.0430 0.0918
   -w 15sec Audio Cond. 0.3676 0.3682 0.3901 0.3093 0.0172 0.0895 1.633 0.0485 0.1008
@@
Table 2: Objective evaluation: Reverb, Dynamics, Amplitude, and Stereo improvement. Clip values are multiplied by 1000.
Category Text relevance Quality1 Quality2 Consistency Preference
EQ 5.11 $\pm$ 0.55 4.55 $\pm$ 0.44 5.01 $\pm$ 0.52 5.05 $\pm$ 0.42 4.84 $\pm$ 0.49
Reverb 5.59 $\pm$ 0.57 4.25 $\pm$ 0.68 5.28 $\pm$ 0.57 5.31 $\pm$ 0.54 5.27 $\pm$ 0.75
Dynamics 4.56 $\pm$ 0.80 4.09 $\pm$ 0.54 5.00 $\pm$ 0.81 5.00 $\pm$ 0.84 4.91 $\pm$ 1.01
Amplitude 6.19 $\pm$ 0.66 3.66 $\pm$ 0.43 5.38 $\pm$ 0.77 5.59 $\pm$ 0.81 5.75 $\pm$ 0.84
Stereo 5.50 $\pm$ 1.05 4.62 $\pm$ 0.83 5.50 $\pm$ 0.63 5.38 $\pm$ 0.70 5.25 $\pm$ 0.77
Mixed degradations 5.28 $\pm$ 0.51 3.85 $\pm$ 0.88 4.60 $\pm$ 0.59 4.65 $\pm$ 0.62 4.93 $\pm$ 0.57
Table 3: Listening study results - Mean Opinion Score with 95% confidence interval.
@@
   -w 5sec Audio Cond. 0.070 0.703 0.624 7.733 0.083 1.175 0.588 7.637 0.075 0.905 0.609 7.692
   -w 15sec Audio Cond. 0.069 0.694 0.623 7.742 0.083 1.161 0.588 7.650 0.073 0.894 0.608 7.702
@@
Table 4: Objective evaluation with FAD, KL, SSIM, and PQ metrics. For readability, KL values were multiplied by 1000.
Training with the rectified flow:
We train SonicMaster using rectified flow (Liu, Gong, and Liu 2022; Esser et al. 2024), to predict the flow velocity from degraded to clean audio in latent space, unlike other models that map noise to output distributions (Fei et al. 2024; Hung et al. 2024).
We assign timestep $t=1$ to the latent representation of the degraded audio $x_{1}$, and $t=0$ to the latent representation of the clean audio target $x_{0}$. During training, we feed the model with samples $x_{t}$, which are interpolations of degraded input $x_{1}$ and clean target $x_{0}$ on a straight line:
$x_{t}=tx_{1}+(1-t)x_{0}$ (1)
where timestep $t$ is drawn from a skewed distribution with increasing probability for higher $t$ as follows:
$p(t)=0.5U(t)+0.5t,\quad t\in[0,1]$ (2)
where $U$ represents a uniform distribution. This skewed distribution gives emphasis to more degraded inputs given the interpolation of training data in Eq. 1.
The model is trained to predict the flow velocity $v_{t}$ from the current $x_{t}$ to the target clean audio $x_{0}$:
$v_{t}=\frac{dx_{t}}{dt}=x_{0}-x_{1}$ (3)
The model $f_{\theta}$ with parameters $\theta$ estimates the velocity $\hat{v_{t}}$:
$f_{\theta}(x_{t},t,c_{text})=\hat{v_{t}}$ (4)
where $c_{text}$ is the text condition from the FLAN-T5 model, which is passed to the dual-stream MM-DiT blocks as one of the streams. The $c_{text}$ condition is also passed through a pooled projection and used to control the scale and shift factors of the adaptive layer-norm layers in both MM-DiT and DiT blocks. The training loss is then given as:
[…]
During inference, the degraded audio input $x_{1}$ is converted into a clean audio output $x_{0}$ by integrating the predicted velocity $\hat{v_{t}}$ using the forward Euler method:
$x_{t+h}=x_{t}+h\hat{v_{t}}$ (6)
where $h\in[0,1]$ is the step computed as the inverse of total timesteps dedicated for integration.
Inference
During inference, SonicMaster takes in an audio input and a text instruction given by user to perform desired restoration/mastering operation. Inference is possible without text input in the so-called auto-correction mode. To process full-length songs, SonicMaster operates on chunks of 30 sec and then connects the segments together. After the first segment is inferred, the last 10 sec of this output are used to condition the next segment inference through the audio pooling branch. The overlapping regions of the resulting segments are then linearly interpolated over the overlapping 10 seconds to connect the segments together.
Experimental Setup and Baselines
Baselines and training setup
We train SonicMaster using NVIDIA L40S GPUs for 40 epochs. We adopt classifier-free guidance (Ho and Salimans 2022) by (i) dropping the text prompt in 1010% of samples and (ii) replacing it in another 1010% with one of four generic phrases (“Make it sound better!”, “Master this track for me, please!”, “Improve this!”, “Can you improve the sound of this song?”). In 2525% of cases, the model is additionally conditioned—via the pooling branch—on the first 10 s of clean audio. Unless stated otherwise, all experiments follow these conditioning settings while comparing multiple SonicMaster variants and baselines.
[…]
Objective Evaluation
Evaluation is conducted along two orthogonal axes. (i) Global perceptual fidelity is quantified with Frechét Audio Distance (FAD) on CLAP embeddings (Elizalde et al. 2023), Kullback–Leibler divergence (KL), structural similarity (SSIM) on 128-bin mel-spectrograms, and the Production Quality (PQ) score from the Audiobox Aesthetics toolbox (Tjandra et al. 2025). (ii) Degradation-specific restoration efficacy is measured by average absolute error reduction: for every degraded clip in a $7000$ clip test set, we compute the relevant (based on the degradation deployed) artefact-aware metric against its clean counterpart from a $1000$ sample reference set, then recompute the metric after SonicMaster processing; the relative decrease indicates how closely each model variant approaches the ground-truth.
For X-band EQ and microphone-TF degradations we compute the spectral balance over eight bands and report their cosine distance. All other EQ effects are scored by the energy ratio between the affected band and the full spectrum. Compression is measured as the standard deviation of frame-level RMS (2048-sample frames, 1024 hop); punch as the mean onset-envelope value (librosa.onset.onset_strength). Because RT60 estimates are unreliable on dense mixes, reverb is assessed via the Euclidean distance of modulation spectra. Clipping uses spectral flatness; volume, the global RMS; and stereo width, the RMS ratio of the mid and side signals, $\mathrm{RMS}\!\left[\frac{L-R}{2}\right]/\mathrm{RMS}\!\left[\frac{L+R}{2}\right]$.
Subjective Evaluation
(a) Ground truth audio.
(b) Degraded by reverb (smearing).
(c) After SonicMaster: reverb removed.
Figure 4: Comparison of spectrograms: (a) ground truth, (b) degraded with reverb, and (c) the output of SonicMaster where smearing is removed. Prompt: ”Please, reduce the strong echo in this song.”
Results
The results of the attribute controllability evaluation are depicted in Tables 1 and 2. SonicMaster significantly outperforms the baselines of Text2FX in EQ, and WPE and HPSS in Reverb. SonicMaster shows a clear improvement in all categories when compared to the degraded and reconstructed inputs. Furthermore, we can observe that the reconstructed input metrics are overall slightly worse (with exceptions) than those of the ground truth degraded inputs.
@@ Results
Across the audio-quality metrics in Table 4, SonicMaster outperforms the degraded inputs in both PQ and KL. Its FAD is only marginally higher than that of the degraded audio, yet markedly lower than the reconstructed baseline. Furthermore, SonicMaster achieves a significant increase in PQ, almost reaching the level of ground truth mastered reference. In SSIM, SonicMaster exhibits lower scores than the degraded inputs but achieves superior performance compared to the reconstruction baseline. From Tables 1, 2, 4, we can observe that SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. SonicMasterMedium (4/12) performs slightly better than the SonicMasterSmall, but still lacks behind SonicMasterLarge in Clip. SonicMasterMedium (6/6) performs the worst out of all variants across all metrics.
As part of our ablation study, we compared SonicMasterLarge with different settings. Models with different lengths of audio conditioning (5 sec, 15 sec) perform similarly to the 10 sec setting. We chose to use 10 sec setting as our default as a compromise between computation speed and length of overlap in neighbouring segments. Further, conditioning during inference, or no conditioning during training and inference at all produced comparable results. While the no condition variant provides best performance in Boom with absolute average error of 0.0658, it also provides the worst performance in Clip error with 2.055 compared to the aforementioned variants. Furthermore, we evaluated inference without text prompts, which yields comparable performance in FAD, SSIM, and PQ, but shows worse performance in KL (0.917 vs 0.696 on single degradation snippets). Performance in Clip and Stereo is also worse with values of 2.812 and 0.1416 as opposed to 1.506 and 0.1058. In EQ, values across the error values across the board are slightly higher for this variant, as the model is not targetting a specific EQ enhancement due to the missing text prompt.
We evaluated Euler solvers with $1$, $10$ (baseline), and $100$ steps, plus a $10$-step 4th order Runge–Kutta (Dormand and Prince 1980) solver. Euler-1 matches the baseline overall but is weaker on Boom, Microphone, Clip, all Reverb subtasks, and shows higher KL. Euler-100 boosts Reverb and Punch yet lowers every EQ score versus the 1-/10-step runs. Runge–Kutta-10 equals Euler-10 on most metrics and tops Clip, but its inference is significantly slower.
When evaluating SonicMaster with full songs, we can again observe significant improvements in EQ metrics (Table 1), and in most functions in Table 2. However, the reverb metric of MSD shows mixed results. In audio quality metrics, SonicMaster shows a decrease in SSIM and FAD compared to degraded inputs, except for FAD in double and triple degradation samples, which points to SonicMaster’s ability to handle multiple degradations at once.
@@
Subjective evaluation results are depicted in Table 3. In text relevance, we can observe very good ratings in Amplitude (6.19) and Reverb categories (5.59), which highlight significant perceptive improvement in declipping, volume increase, and in dereverberation. These two categories further show the highest improvement in quality, highest consistency ratings, and highest preference. EQ shows the third best text relevance, but worst preference ratings. Overall, SonicMaster samples are rated higher in quality compared to inputs and preferred across the board. A paired t-test on Quality1 and Quality2 ratings shows statistically significant differences in Reverb, Amplitude, Stereo, and mixed degradations. These results further confirm SonicMaster’s capabilities in text-controllable music restoration and mastering.
Discussion
The experiments confirm that SonicMaster improves audio quality, proving that a generative strategy for music restoration and mastering is feasible when backed by a large corpus and a suitable training objective. A key limitation, however, are occasional artefacts introduced by the lossy latent representation: reconstructed clips can exhibit “robotic” vocals or muted instruments, especially for certain genres, and the effect is stronger for degraded inputs than for their clean references. Future research should pursue less lossy latent encoders or alternative representations to minimize these distortions.
The observed decrease in SonicMaster’s performance on full songs in SSIM and Reverb metrics could be related to the way neighbouring segments are connected together. Improving on this aspect could increase the objective performance further. Moreover, evaluating the amount of reverberation in music is challenging, as the audio is usually very dense. Additionally, the way a generative model such as SonicMaster removes reverb in the latent space is not telegraphed, which makes it more difficult to choose a correct metric. A deep exploration of this topic would be of use to the community.
Conclusion
We introduced SonicMaster, the first unified text-guided generative model for music restoration and mastering, capable of addressing a wide range of audio degradations within a single model. Key contributions include the creation of a new SonicMaster dataset [^2] Dataset, code, samples, model, and demo available through https://amaai-lab.github.io/SonicMaster/ of paired degraded and high-quality music with textual annotations and the use of a flow-matching training paradigm to directly learn the restoration mapping. We also developed a text-conditioned multimodal generative architecture that unifies diverse enhancement tasks, enabling the model to perform all restoration and mastering steps in one generative pass. Comprehensive evaluations demonstrate that SonicMaster substantially improves audio quality, achieving strong results in objective metrics (e.g., FAD, KL divergence, SSIM, PQ) and consistently earning higher listener preferences in blind tests compared to both original and baseline outputs.

### v2

@@ 4.1 Baselines and training setup
   -w 5s Audio Cond. 0.0111 0.0716 0.0021 0.0061 0.0058 0.0386 0.0605 0.0628 0.0124 0.0387
   -w 15s Audio Cond. 0.0117 0.0750 0.0020 0.0064 0.0063 0.0320 0.0552 0.0525 0.0079 0.0398
@@
Table 1: EQ Objective evaluation (average absolute error) – the lower, the better.
Evaluation is conducted along two orthogonal axes. (i) Global perceptual fidelity is quantified with FAD on CLAP embeddings (Elizalde et al., 2023), Kullback–Leibler divergence (KL), structural similarity (SSIM) on 128-bin mel-spectrograms, and the Production Quality (PQ) score from the Audiobox Aesthetics toolbox (Tjandra et al., 2025). (ii) Degradation-specific restoration efficacy is measured by average absolute error reduction: for every degraded clip in a $7000$ clip test set, we compute the relevant (based on the degradation deployed) artefact-aware metric against its clean counterpart from a $1000$ sample reference set, then recompute the metric after SonicMaster processing; the relative decrease indicates how closely each model variant approaches the ground-truth.
For X-band EQ and microphone-TF degradations, we compute the spectral balance over eight bands and report their cosine distance. All other EQ effects are scored by the energy ratio between the affected band and the full spectrum. Compression is measured as the standard deviation of frame-level RMS (2048-sample frames, 1024 hop); punch as the mean onset-envelope value (librosa.onset.onset_strength). Because RT60 estimates are unreliable on dense mixes, reverb is assessed via the Euclidean distance of modulation spectra. Clipping uses spectral flatness; volume, the global RMS; and stereo width, the RMS ratio of the mid and side signals, $\mathrm{RMS}\!\left[\frac{L-R}{2}\right]/\mathrm{RMS}\!\left[\frac{L+R}{2}\right]$.
Furthermore, to benchmark against existing methods, we conducted an additional study with 13 participants comparing SonicMaster against Text2FX (Chu et al., 2025) and Mel2Mel + Diffwave (Kandpal et al., 2022) on 20 randomly selected samples from our test set. The evaluation included 10 samples with X-band EQ degradation and 10 with reverberation artifacts. Note that Text2FX is limited to EQ effects (as its reverb effect is only additive, thus excluded), making direct comparison feasible only for the EQ category. Since the baseline methods’ evaluation sets are not publicly available, we performed this comparison exclusively on our curated test data.
5 Results
5.1 Objective Evaluation
Degradation-Specific Performance: Tables 4.1 and 5.1 demonstrates SonicMaster’s superiority over baselines of Text2FX in EQ, and WPE/HPSS in Reverb. SonicMaster improves in all categories when compared to the degraded and reconstructed inputs. Furthermore, the reconstructed input metrics are overall slightly worse (with exceptions) than those of the ground truth degraded inputs.
@@ 5.1 Objective Evaluation
   -w 5s Audio Cond. 0.3717 0.3658 0.3919 0.3079 0.0164 0.0893 1.779 0.0430 0.0918
   -w 15s Audio Cond. 0.3676 0.3682 0.3901 0.3093 0.0172 0.0895 1.633 0.0485 0.1008
@@
Table 2: Objective scores: Reverb, Dynamics, Amplitude, and Stereo. Clip are multiplied by 1000.
@@
   -w 5s Audio Cond. 0.070 0.703 0.624 7.733 0.083 1.175 0.588 7.637 0.075 0.905 0.609 7.692
   -w 15s Audio Cond. 0.069 0.694 0.623 7.742 0.083 1.161 0.588 7.650 0.073 0.894 0.608 7.702
@@
Table 3: Objective scores: FAD, KL, SSIM, and PQ. For readability, KL are multiplied by 1000.
@@
Perceptual Quality Assessment: Table 5.1 reveals SonicMaster outperforms the degraded inputs in both PQ and KL. FAD is marginally higher than that of the degraded audio, yet markedly lower than the reconstructed baseline. Furthermore, SonicMaster achieves a significant increase in PQ, almost reaching the level of ground truth mastered reference. In SSIM, SonicMaster exhibits lower scores than degraded inputs but achieves superior performance compared to the reconstruction baseline.
Architecture Scaling Analysis: Tables 4.1, 5.1, 5.1, reveal interesting scaling dynamics. SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. SonicMasterMedium (4MM-DiT/12DiT) performs slightly better than the SonicMasterSmall, but still lacks behind SonicMasterLarge in Clip. SonicMasterMedium (6MM-DiT/6DiT) performs the worst out of all variants across all metrics.
5.2 Ablation Studies
Audio Conditioning Duration: We evaluated SonicMasterLarge with different conditioning lengths (5s, 10s, 15s), finding comparable performance across configurations. The 10-second setting balances computational efficiency with temporal overlap for long-form processing.
Conditioning Strategy Analysis: The no-conditioning variant achieves optimal Boom correction (0.0658 absolute error) but poor Clip restoration (2.055 vs. standard variants), highlighting multi-modal guidance importance for challenging tasks.
Text Prompt Dependency:Inference without text prompts maintains comparable FAD, SSIM, and PQ but shows degraded KL divergence (0.917 vs. 0.696). Critical drops occur in Clip restoration (2.812 vs. 1.506) and Stereo processing (0.1416 vs. 0.1058), with elevated EQ errors. This confirms text conditioning enables targeted restoration rather than generic improvements.
Method CE $\uparrow$ CU $\uparrow$ PC $\uparrow$ PQ $\uparrow$
Original $6.94_{\pm 0.48}$ $7.29_{\pm 0.43}$ $3.45_{\pm 0.36}$ $6.70_{\pm 0.50}$
LTAS-EQ $6.77_{\pm 0.54}$ $7.04_{\pm 0.57}$ $3.75_{\pm 0.45}$ $6.49_{\pm 0.57}$
BEHM-GAN $6.82_{\pm 0.43}$ $7.19_{\pm 0.44}$ $3.47_{\pm 0.35}$ $6.63_{\pm 0.56}$
BABE $\textbf{6.96}_{\pm 0.37}$ $\textbf{7.32}_{\pm 0.37}$ $3.32_{\pm 0.29}$ $6.79_{\pm 0.36}$
BABE-2 $6.79_{\pm 0.34}$ $7.16_{\pm 0.29}$ $3.46_{\pm 0.28}$ $\textbf{7.05}_{\pm 0.27}$
SonicMaster $6.87_{\pm 0.55}$ $7.25_{\pm 0.50}$ $\textbf{3.86}_{\pm 0.39}$ $6.93_{\pm 0.52}$
Table 4: Comparison of mean across metrics CE, CU, PC, and PQ.
Long-Form Audio Evaluation: Full-song evaluations confirm SonicMaster’s effectiveness, with substantial improvements in EQ-related metrics (Table 4.1) and most degradation functions (Table 5.1). Reverberation metrics show mixed results, likely due to increased complexity of spatial processing in extended musical contexts where room acoustics interact with diverse instrumental timbres and dynamic variations. SSIM and FAD decrease compared to degraded inputs, except for FAD in multi-degradation samples, indicating SonicMaster’s ability to handle compound degradations.
5.3 Piano Recordings Evaluation
To test SonicMaster generalization, we evaluate historical solo piano pieces [^3] http://research.spa.aalto.fi/publications/papers/dafx-babe2/ using established baselines: LTAS-EQ, BEHM-GAN (Moliner & Välimäki, 2023) model for bandwidth extension, and BABE/BABE-2 diffusion-based generative equalizers (Moliner et al., 2024; Moliner & Välimäki, 2023). BABE-2 represents a state-of-the-art specialized method for old recordings, it uses a diffusion prior to restore lost high frequencies and remove coloration, and has shown impressive improvements in archival music (Moliner et al., 2024). Despite lacking domain-specific training, SonicMaster came surprisingly close to these specialized baselines. In objective evaluations, SonicMaster restored samples achieved a PQ of $6.93$, nearly matching the $7.05$ obtained by BABE-2.
Category Text relevance Quality1 Quality2 Consistency Preference
EQ $5.11_{\pm 0.55}$ $4.55_{\pm 0.44}$ $1_{\pm 0.52} 5.05±0$$5_{\pm 0.42}$ $4.84_{\pm 0.49}$
Reverb $5.59_{\pm 0.57}$ $4.25_{\pm 0.68}$ $5.28_{\pm 0.57}$ $5.31_{\pm 0.54}$ $5.27_{\pm 0.75}$
Dynamics $4.56_{\pm 0.80}$ $4.09_{\pm 0.54}$ $5.00_{\pm 0.81}$ $5.00_{\pm 0.84}$ $4.91_{\pm 1.01}$
Amplitude $6.19_{\pm 0.66}$ $3.66_{\pm 0.43}$ $38_{\pm 0.77} 5.5$$59_{\pm 0.81}$ $5.75_{\pm 0.84}$
Stereo $5.50_{\pm 1.05}$ $4.62_{\pm 0.83}$ $5.50_{\pm 0.63}$ $5.38_{\pm 0.70}$ $5.25_{\pm 0.77}$
Mixed degradations $5.28_{\pm 0.51}$ $3.85_{\pm 0.88}$ $4.60_{\pm 0.59}$ $4.65_{\pm 0.62}$ $4.93_{\pm 0.57}$
Table 5: Listening study - SonicMaster’s performance on specific degradations – MOS 95% CI.
5.4 Subjective Evaluation
@@
Table 5 shows results of the first subjective evaluation test. Text relevance ratings are highest in the Amplitude (6.19) and Reverb categories (5.59), indicating effective declipping, volume increase, and dereverberation. These two categories also show the highest improvement in quality, as well as the highest consistency and preference ratings. EQ shows the fourth-best text relevance, but the worst preference ratings. Overall, SonicMaster samples are rated higher in quality compared to inputs and preferred across the board. A paired $t$-test on Quality1 and Quality2 ratings shows statistically significant differences in Reverb, Amplitude, Stereo, and mixed degradations.
Category SonicMaster Mel2Mel + Diffwave (Kandpal et al., 2022) Text2FX (Chu et al., 2025)
Reverberation 128/130 (98.5%) 2/130 (1.5%)
EQ (X-band) 122/130 (93.8%) 6/130 (4.6%) 2/130 (1.5%)
Table 6: Listening study - comparative evaluation results - 13 participants × 10 samples.
The comparative evaluation against existing baselines demonstrates SonicMaster’s superior performance across both reverb and EQ degradation categories (Table 6). For reverb artifacts, participants overwhelmingly preferred SonicMaster over Mel2Mel + Diffwave (Kandpal et al., 2022), selecting our method in 128 out of 130 total comparisons (10 samples × 13 participants), with Mel2Mel + Diffwave chosen only twice. In the EQ category, SonicMaster achieved similarly strong results with 122 out of 130 preferences, while Mel2Mel + Diffwave received 6 votes and Text2FX (Chu et al., 2025) garnered 2 votes. These results underscore SonicMaster’s effectiveness in addressing both spatial acoustic degradations and spectral imbalances compared to specialized baseline methods.
6 Discussion
Experiments confirm that SonicMaster’s generative approach is effective when trained on a large corpus with a suitable objective. The historical piano experiment demonstrated SonicMaster’s strong generalization: even on out-of-domain, severely degraded audio, it produced enhancements close to the best specialized solution, BABE-2. This highlights the potential of general-purpose audio restoration AI. However, a key limitation is that the lossy latent representation can introduce artifacts, such as robotic vocals or muted instruments, especially in certain genres. The observed decrease in SonicMaster’s performance on full songs in SSIM and Reverb metrics could be related to the way neighbouring segments are connected together. Improving on this aspect could increase the objective performance further. Evaluating reverberation in dense music is challenging, and how SonicMaster removes it in latent space is not explicitly observable, making metric selection difficult. A deeper study of this issue would benefit the community.
7 Conclusion
We introduced SonicMaster, the first unified text-guided generative model for music restoration and mastering, capable of handling 19 diverse degradations within a single framework. Our contributions further include the creation of a paired degraded–clean dataset with textual annotations, the introduction of a flow-matching paradigm for directly learning restoration mappings, and the integration of natural language conditioning for precise and flexible control. Evaluations show that SonicMaster consistently improves the audio quality, outperforming baselines in terms of objective metrics and listener studies. It also achieved strong zero-shot performance on old piano recordings, highlighting its versatility suggesting a path toward a generalist restoration framework–one capable of addressing diverse challenges through prompt guidance while approaching the quality of specialist methods.
Acknowledgements
This work has received support from SUTD’s Kickstart Initiative under grant number SKI 2021 04 06, MOE Tier 2 under grant number MOE-T2EP20124-0014, and SMU-SUTD grant number 2023_02_04.